OpenStudy (anonymous):

One number is 5 more than another number. What is the smaller number if their product is -3?

OpenStudy (anonymous):

Let the two numbers be x and y. It is given that x = 5+y and xy = -3 => (5+y)*y =-3 5y+y^2 =-3 y^2+5y +3 =0 I am stuck here. Can someone solve the equation?